# Adaptive Querying for Reward Learning from Human Feedback

### v1

@@ V-A Results and Discussion
Method Navigation: unavoidable NSE Vase: unavoidable NSE Boxpushing: avoidable NSE Freeway: avoidable NSE
Oracle $51.37\pm 2.69$ $54.46\pm 6.70$ $44.62\pm 9.97$ $3759.8\pm 0.0$
Naive $36.11\pm 1.39$ $36.0\pm 2.89$ $39.82\pm 5.44$ $61661.0\pm 0.0$
RI $40.10\pm 0.69$ $37.42\pm 1.01$ $42.15\pm 2.44$ $71716.6\pm 0.0$
Ours $49.18\pm 13.67$ $63.0\pm 0.73$ $46.17\pm 0.86$ $1726.5\pm 0.0$
Effect of learning using AFS We first examine the benefit of querying using AFS, by comparing the resulting average NSE penalties and the cost for task completion, across domains and query budget. Figure 5 shows the average NSE penalties when operating based on an NSE model learned using different querying approaches. Clusters for critical state selection were generated using KMeans clustering algorithm with $K\!=\!3$ for navigation, vase and safety-gym’s push domains (Figure 6 (a-c)) and $K\!=\!5$ for the Atari Freeway domain (Figure 6 (d)). The results show that our approach consistently performs similar to or better than the baselines.
There is a trade-off between optimizing task completion and mitigating NSEs, especially when NSEs are unavoidable. While some techniques are better at mitigating NSEs, they significantly impact task performance. Table I shows the average cost for task completion. Lower values are better for both NSEs and task completion cost. While the Naive Agent has a lower cost for task completion, it incurs the highest NSE penalty as it has no knowledge of $R_{N}$. RI causes more NSEs, especially when they are unavoidable, as its reward function does not fully model the penalties for mild and severe NSEs. Overall, the results show that our approach consistently mitigates avoidable and unavoidable NSEs, without affecting the task performance substantially.
(a) Navigation: Unavoidable NSE
(b) Vase: Unavoidable NSE
(c) Safety-gym Push
(d) ATARI Freeway
Effect of learning from multiple feedback formats To better understand the need for and benefits of AFS, we investigate the benefits of learning from more than one feedback type in general. Figure 6 shows the results comparing the average NSE penalties of learning from a single feedback and multiple feedback formats, with varying budget for querying across domains. In the single feedback case(Figure 6(a-c)), Corrections format successfully mitigates NSEs with fewer feedback across domains. However, its reliance on constant human guidance is a limitation. While Demo-Action Mismatch requires less human guidance, it is less effective in avoiding NSEs. The effectiveness of Demo-Action Mismatch improves significantly depending on its position within a sequence of feedback formats (Figure 6(d-f)). For instance, using Demo-Action Mismatch before Corrections, in safety-gym’s push domain, results in a lower average penalty with a smaller budget. However, in the vase domain with unavoidable NSEs, the agent performs better when Demo-Action Mismatch follows Corrections. On the other hand, Approval and Annotated Approval have a similar performance across domains and require higher samples to learn the true distribution of NSEs. However, when combined with Corrections or Annotated Corrections, the performance improves considerably (Figure 6(g-i)). Learning the underlying NSE severities demands a significantly higher number of samples when using a combination of Approval and Annotated Approval formats. Likewise, while Ranking is less effective in mitigating NSEs, it is more effective when used in combination with Corrections or Annotated Corrections (Figure 6(j-l)).
These results show that learning from more than one feedback format is generally useful but the benefits depend on the formats considered together and the order in which they are combined. Identifying the right ordering of feedback formats a priori is often practically infeasible. Our approach enables the agent to identify when and how to query, without any additional pre-processing.
[…]
The results in Figure 7 demonstrate that increasing $K$ generally improves the performance of our approach, with both clustering methods. A higher number of clusters allows for a more refined grouping of states based on distinct state features, enabling the agent to query the human for feedback across a more diverse set of states. This diversity enhances the agent’s ability to accurately learn and mitigate NSEs.
VI Conclusion
@@
The proposed Adaptive Feedback Selection (AFS) facilitates querying a human in different formats in different regions of the state space, to effectively learn a reward function. Our approach uses information gain to identify critical states for querying, and the most informative feedback format to query in these states, while accounting for the cost and uncertainty of receiving feedback in each format. Our empirical evaluations using four domains in simulation demonstrate the effectiveness and sample efficiency of our approach in mitigating avoidable and unavoidable negative side effects (NSEs), based on explicit and implicit feedback formats. In the future, we aim to validate our assumptions and results using user studies and extend our approach to continuous settings.

### v2

@@ V-A Results and Discussion
VI In-Person User Study with a Physical Robot Arm
We conducted an in-person study with a Kinova Gen3 7DoF arm [25] tasked with delivering two objects—an orange toy and a white box—across a workspace containing items of varying fragility (Figure 9). This setup involves users providing both interface-based and kinesthestic feedback to the robot. The study was approved by Oregon State University IRB. Participants were compensated with a $\mathdollar 15$ Amazon gift card for their participation in the study.
This user study had three goals: (1) to measure our approach’s effectiveness in reducing NSEs for a real-world task, (2) to understand how users perceive the adaptivity, workload and competence of the robot operating in the AFS framework, and (3) to evaluate the extent to which AFS captures user preferences in practice, while ensuring maximum information gain during the learning process.
VI-A Methods
VI-A1 Participants
We conducted a pilot study in simulation to inform our overall design, the details of which are discussed in Appendix B. We conducted another pilot study with N=10N\!=\!10 participants to evaluate the study setup with the Kinova arm. In particular, this pilot study assessed the clarity of instructions, survey wording, and feasibility of the task design in the object delivery task of the Kinova arm. Based on the participant feedback, we simplified the survey questions and included example trajectories that demonstrated safe and NSE-causing behaviors. For the main study, we recruited N=30N\!=\!30 participants with basic computer literacy from the general population through university mailing lists and public forums. Participants were aged 18187272 years (M=32.10,SD=13.11M\!=\!32.10,SD\!=\!13.11), with $53.3\%$ men and $46.7\%$ women. Participants reported varied prior experience with robots: $73.3\%$ had general awareness of similar robot products, $6.7\%$ had researched or investigated robots, $3.3\%$ had interacted through product demos, and $13.3\%$ had no prior awareness of similar products.
VI-A2 Robotic System Setup
The Kinova Gen3 arm was equipped with a joint space compliant controller which allowed participants to physically move the joints of the arm through space with gravity compensation when needed. Additionally, a task-space planner allowed for navigation to discrete grid positions for both feedback queries and policy execution [25]. Figure 9(a) shows the physical workspace and the two delivery objects, while Figure 9(b) shows the corresponding PyBullet simulation used for visualization during GUI-based feedback. A dialog box was displayed to prompt the participant whenever feedback was queried [^1] See Appendix C for details on the dialog box and examples for each feedback format..
VI-A3 Interaction Premise
The interaction simulated an assistive robot delivering objects to their designated bins. Specifically, the task required the Kinova arm to deliver an orange plush toy and a rigid white box to their respective bins while avoiding collision with surrounding obstacles of different fragility. Collisions with fragile obstacles (e.g. a glass vase) during delivery of the plush toy were considered a mild NSE. Collisions involving the white rigid box were severe NSEs if with a fragile object and were mild NSEs if with a non-fragile object. All other scenarios were considered safe. The workspace was discretized into a grid of cells marked with tape on the tabletop and mirrored in the GUI. Each cell represented a state corresponding to possible end-effector position.
Figure 9: Task setup for the human subject study. (a) Physical setup of the task for human subjects study; (b) Replication of the physical setup using PyBullet. A dialog box corresponding to the current feedback format is shown for every query.
VI-A4 Study Design
The robot’s state space was discretized and represented as $\langle x,y,i_{1},i_{2},o,f,g_{1},g_{2}\rangle$, where $(x,y)$ denote the end-effector position, $i_{1}$ and $i_{2}$ indicate the presence of either orange plush toy or white rigid box in the end effector, $o$ indicates the presence of an obstacle, and $f$ indicates obstacle fragility, and $g_{1}$ and $g_{2}$ indicate whether either of the objects were delivered in their corresponding goal locations (i.e., orange plush toy in white bin and the white box in the wicker bin).
Participants interacted with the robot through four feedback formats, $\mathcal{F}=\{\text{App, Corr, Rank, DAM}\}$, during both the training and main experience phases. Depending on the feedback format, the Kinova arm executed the queried action in the physical workspace or displayed a simulation of the action in the graphical user interface (GUI). Interaction across the four feedback formats are described below.
  1. 1.
    Approval: The robot executed a single action in simulation, and participants indicated whether it was safe by selecting “yes” or “no” in the GUI.
  2. 2.
    Correction: The robot first executes action prescribed by its policy in simulation. If the action in simulation is deemed unsafe by the participant, the robot in the physical setup moves to the queried location. Participants then correct the robot by physically moving the robot arm to demonstrate a safe alternative action.
  3. 3.
    Demo-Action Mismatch: The robot first physically moved its arm to a specific end-effector position in the workspace. Participants then provided feedback by guiding the arm to a safe position, thereby demonstrating the safe action. The robot compares the action given by its policy to the demonstrated action. If the robot’s action and the demonstrated actions do not match, then the robot’s action is considered unsafe.
  4. 4.
    Ranking: Simulation clips of two actions selected at random in a given state were presented in GUI. Participants compared the two candidate actions and selected which was safer. If both actions were judged equally safe or unsafe, either option could be chosen.
Each participant experienced four learning conditions in a within-subjects, counterbalanced design:
  1. 1.
    The baseline RI approach proposed in [16],
  2. 2.
    AFS with random $\Omega$, where critical states are randomly selected,
  3. 3.
    AFS with a fixed feedback format (DAM) for querying, consistent with prior works that rely primarily on demonstrations, and
  4. 4.
    The proposed AFS approach, where both the feedback format and the critical states are selected to maximize information gain.
Each condition is a distinct feedback query selection strategy controlling how the robot queried participants during learning. These conditions are the independent variables. The dependent measures include NSE occurrences, their severity, perceived workload, trust, competence and user alignment.
VI-A5 Hypotheses
We test the following hypotheses in the in-person study. These hypotheses were derived from trends observed in the experiments and human subjects study in simulation (Sec.  V and Appendix B).
H1: Robots learning using AFS will have fewer NSEs in comparison to the baselines.
This hypothesis is derived from the results of our experiments on simulated domains (Figure 8) where AFS consistently reduced NSEs while completing the assigned task. We hypothesize that this trend extends to physical human-robot interactions.
H2: AFS will achieve comparable or better performance compared to the baselines, with a lower perceived workload for the users.
The results on simulated domains (Figure 7) show that AFS achieved better or comparable performance to the baselines, using fewer feedback queries. While the in-person user study requires relatively greater physical and cognitive effort, we expect the advantage of the sample efficiency to persist and investigate whether it translates to reduced perceived workload.
H3: Participants will report AFS as more trustworthy, competent, and aligned with user expectations, in comparison to the baselines.
In the human subjects simulation study (Table III in the Appendix), participants reported that AFS selected intelligent queries, targeted critical states, and improved the agent’s performance, reflecting indicators of trust, competence and user alignment. We hypothesize that this trend extends to physical settings as well.
Hypotheses H1 and H2 explore trends identified in simulation and are therefore confirmatory. Hypothesis H3 builds on the perception measures used in the human subjects study in simulation, and is hence treated as an extended confirmatory hypothesis.
VI-A6 Procedure
Each study session lasted approximately one hour and followed three phases.
Training
Participants were first introduced to the task objective, workspace, and the four feedback formats. For each format, they provided feedback on four sample queries to practice both GUI-based and kinesthetic interactions. After the completing each format, the participants rated the following: (i) probability of responding to a query in that format, $\psi(f)$, (ii) perceived cost or effort required to provide feedback, $C(f)$, and (iii) the overall task workload. This phase helped establish measures like feedback likelihood, perceived effort, and workload.
Main Experience
Following training, participants completed the four learning conditions corresponding to different approaches under evaluation. In each condition, the participants provided feedback to train the robot to avoid collision while performing the object-delivery task. Depending on the feedback format selected by the querying strategy, participants either evaluated short simulation clips on the GUI or physically guided the robotic arm. At the end of each condition, the robot executed its learned policy based on its learning under that condition. The participants then observed its performance and completed a brief post-condition questionnaire assessing workload, trust, perceived competence, and user-alignment.
Closing
At the end of the study, participants compared the four learning approaches in terms of trade-offs between learning speed and safety. Participants reported their preferences on providing feedback through multiple formats versus relying on a single feedback format. These responses offered qualitative insight into AFS’s practicality and user acceptance.
Figure 10: Results from the user study on the Kinova 7DoF arm.
VI-A7 Measures
We collected both quantitative and qualitative measures. The quantitative measure captured task-level performance through the frequency and the severity of NSEs (mild and severe). Qualitative measures captured participants’ perceptions of the following.
  1. 1.
    Workload: Participants’ perceived workload across the feedback formats and learning conditions were measured using the NASA Task Load Index (NASA TLX) [18]. The questionnaire scales were transformed to 7-point subscales ranging from “Very Low” (1) to “Very High” (7). Responses were collected during the training phase and after each condition in the main experience phase.
  2. 2.
    Robot Attributes: Perceived robot attributes, like competence, warmth and discomfort, were measured using the 9-point Robotic Social Attributes Scale (RoSAS) [10], ranging from “Strongly Disagree” (1) to “Strongly Agree” (9). Participants completed this questionnaire after each learning condition.
  3. 3.
    Trust: A custom 10-point trust scale $(0\%-100\%)$ was used to measure participants’ confidence in the robot’s ability to act safely under each learning condition. Participants rated their trust both before and after the robot’s training phase to capture changes in its learning performance.
  4. 4.
    User Alignment: Participants’ perception of user alignment was assessed using a custom 7-point Likert scale ranging from “Strongly Disagree” (1) to “Strong Agree” (7). Participants rated (i) how well the critical states queried by the robot aligned with their own assessment of which states were important for learning, and (ii) how well the feedback formats chosen across conditions matched their personal feedback preferences. Higher rating indicated stronger perceived alignment between the robot’s querying strategy and the participants’ expectations.
VI-A8 Analysis
Survey responses were compiled into cumulative RoSAS (competence, warmth, discomfort) and NASA-TLX workload scores. A repeated-measures ANOVA (rANOVA) tested for significant differences across learning conditions; we report the $F$-statistic, $p$-value and effect size as generalized eta-squared ($\eta^{2}_{G}$). When effects were significant, Tukey’s post-hoc tests identified pairwise differences. All results are reported with means (M), standard errors (SE), and $p$-values.
VI-B Results
We evaluate hypotheses H1-H3 using both objective and subjective measures. Data from all 30 participants were included in the analysis, as all sessions were completed successfully.
VI-B1 Effectiveness of AFS in Mitigating NSEs (Hypothesis H1)
Figure 10(a) shows the average penalty incurred under each condition. AFS approach incurred the least NSE penalty ($M=3.83,SE=1.21$), substantially lower than AFS with random $\Omega$ ($M=11.55,SE=1.57$) and AFS with a fixed feedback format ($M=10.50,SE=0.37$). The RI baseline incurred higher penalties ($M=5.00,SE=0.00$) compared to AFS. These results confirm hypothesis H1 and demonstrate that adaptively selecting both critical states and feedback formats reduced unsafe behaviors more effectively than random or fixed querying strategies.
Figure 11: User study results. (a)-(b) RoSAS competence and NASA Task-Load across the four conditions in the main study; (c) NASA Task-Load across feedback formats.
VI-B2 Learning Efficiency and Workload (Hypothesis H2)
We first compare the perceived workload across different feedback formats, followed by the results across learning conditions. Demonstration is the most widely used feedback format in existing works but was perceived as the most demanding (Figure 11(c)). While corrections offer corrective action in addition to disapproving agent’s action, it also imposed substantial effort on the users. Approval required the least workload but conveyed limited information. A repeated-measures ANOVA revealed a significant effect of feedback format on perceived workload, ($F(3,87)=3.33,p=0.023,\eta_{G}^{2}=0.046$). Post hoc comparisons indicated that Approval ($M=2.11,SE=0.12$) imposed significantly lower workload ($p=0.026$) than Demo-Action Mismatch ($M=2.62,SE=0.19$), while no other pairwise differences reached significance. This trade-off underscores the need for an adaptive selection strategy to balance informativeness with user effort.
The rANOVA analysis across the four learning conditions further revealed a significant effect in the NASA-TLX workload ratings ($F(3,87)=3.73,p=0.014,\eta_{G}^{2}=0.030$). Among the four conditions, AFS achieved one of the lowest perceived workload ratings ($M=2.34,SE=0.12$), comparable to AFS with random $\Omega$ ($M=2.26,SE=0.15$) and lower than both AFS with fixed format ($M=2.56,SE=0.19$) and RI ($M=2.64,SE=0.19$). Tukey post-hoc tests showed that workload in AFS with random $\Omega$ imposed a significantly lower workload than RI ($p=0.033$). Overall, these results support H2, indicating that adaptively selecting queries helps reduce perceived workload relative to the baselines (Figure 10(b)).
VI-B3 Trust, Competence, and Preference Alignment (Hypothesis H3)
Participants’ rating on the robot’s ability to act safely increased after learning with AFS, as shown in Figure 11(b). A significant effect was also found for perceived robot competence ($F(3,87)=10.6,p<0.001,\eta_{G}^{2}=0.082$) (Figure 11(a)). AFS was rated highest ($M=7.04,SE=0.32$), significantly greater than AFS with random $\Omega$ ($M=5.88,SE=0.32,p=0.002$) and AFS with fixed format ($M=5.88,SE=0.30,p<0.001$), while comparable to RI ($M=6.68,SE=0.32$). These results support H3—AFS was perceived as more competent and trustworthy compared to the baselines.
Descriptive analyses of user alignment on state criticality and feedback alignment ratings, indicated consistent trends across participants. While differences between conditions were not statistically significant ($p>0.05$), AFS consistently received higher ratings for feedback alignment ($M=3.79,SE=0.42$) relative to state criticality ($M=3.14,SE=0.40$), suggesting that participants found AFS’s query selections relevant and aligned with their preferences. Participants (both those aware and unaware of similar robotic systems) perceived AFS’s queries as critical for learning and well-aligned with their feedback preferences. Participants with prior research experience rated state criticality and format alignment comparable, indicating confidence in adaptivity of AFS’s querying process.
VII Discussion
Our experiments followed an increasingly realistic progression in design. In the experiments in simulation with both avoidable and unavoidable NSEs, AFS incurred lower penalties and overall costs compared to the baselines, demonstrating its ability to balance task performance with safety. The results of our pilot study, where users interacted with a simulated agent, showed that AFS effectively learns the participant’s feedback preference model and uses them to select formats aligned with user expectations. Finally, the in-person user study with the Kinova arm, showed the practicality of using AFS in real-world settings, achieving favorable ratings on trust, workload, and user-preference alignment. These finding support our three hypotheses regarding the performance of AFS: (H1) it reduces unsafe behaviors more effectively than the baselines, (H2) it improves learning efficiency while reducing user workload, and (H3) it is perceived as more trustworthy and competent. The results collectively highlight that adaptively selecting both the query format and the states to pose the queries to the user enhances learning efficiency and reduces user effort.
Beyond confirming these hypotheses, the findings provide important design implications for human-in-the-loop learning systems. By modeling the trade-off between informativeness and effort, AFS offers a framework to balance user workload with the need for high-quality feedback. The learned feedback preference model allows the agent to adaptively select querying formats while minimizing human effort. Using KL-divergence as stopping criterion further enables adaptive termination of the querying process. This overcomes the problem of determining the “right” querying budget for a problem, and shows that AFS enables efficient learning while minimizing redundant human feedback. These design principles can inform the development of interactive systems that adapt query format and frequency based on agent’s current knowledge and user feedback preferences. Overall the results show that AFS (1) consistently outperforms the baselines across different evaluation settings, and (2) can be effectively deployed in real-world human-robot interaction scenarios.
A key strength of this work lies in its extensive evaluation, from simulation to real robot studies, supporting AFS’s robustness and practicality. One limitation, however, is that the current evaluation focuses on discrete environments. Extending AFS to continuous domains introduces challenges such as identifying critical states and estimating divergence-based information gain in high-dimensional spaces. While gathering feedback at the trajectory-level is relatively easier in continuous settings, gathering state-level feedback, which is the focus of this work, is challenging. These challenges stem from the need for scalable state representations and efficient sampling strategies, which will be a focus for future work.
VIII Conclusion and Future Work
@@
The proposed Adaptive Feedback Selection (AFS) facilitates querying a human in different formats in different regions of the state space, to effectively learn a reward function. Our approach uses information gain to identify critical states for querying, and the most informative feedback format to query in these states, while accounting for the cost and uncertainty of receiving feedback in each format. Our empirical evaluations using four domains in simulation and a human subjects study in simulation demonstrate the effectiveness and sample efficiency of our approach in mitigating avoidable and unavoidable negative side effects (NSEs). The subsequent in-person user study with a Kinova Gen3 7DoF arm further validates these finding, showing that AFS not only improves NSE avoidance but also enhances user trust, competence perception, and user-alignment. While AFS assumes that human feedback reflects a true underlying notion of safety, biased feedback can misguide the robot and lead to unintended NSEs. Understanding when such biases arise and how to correct for them remains an open challenge. Extending AFS with bias-aware inference mechanisms is a promising future direction. Future work will also focus on extending AFS to continuous state and action spaces, strengthening AFS’s applicability to complex, safety-critical domains where user-aware interaction is essential.
@@ ACKNOWLEDGMENT
Appendix A Additional Experiments in simulation
Figure 13: An instance of the Freeway domain, and the average penalty incurred.
A-A Effect of Clustering
To understand how clustering influences the effectiveness of our approach (AFS), we evaluate AFS under different clustering configurations. In particular, we vary the number of clusters and the clustering approach.
[…]
A-B Learning with Implicit and Explicit Feedback Formats
To demonstrate AFS’s flexibility across feedback modalities, we extend it to handle both explicit (e.g., DAM) and implicit (e.g., gaze) feedback formats.
A-B1 Implicit Feedback Format
We consider a human’s gaze on the screen as the implicit feedback format. Here, the robot requests to collect gaze data of the user and compares its action outcomes with the gaze positions of the user [39]. Actions with outcomes aligning with the average gaze direction are labeled as acceptable ($l_{a}$), and unacceptable ($l_{h}$) otherwise.
A-B2 Domain
We evaluate AFS in the Atari Freeway environment, where the robot (a chicken) navigates ten cars moving at varying speeds to reach the destination quickly while avoiding being hit(Figure 13(a)). Being hit by a car moves the robot back to its previous position, and is a severe NSE. A game state is defined by coordinates $(x_{1},y_{1})$ and $(x_{2},y_{2})$, i.e., the top left and bottom right corners of the robot and cars, extracted from the Atari-HEAD dataset [51]. Similar to [39], only car coordinates within a specific range of the robot are considered. The robot can move up, down or stay in place, with unit cost and deterministic transitions.
A-B3 Effect of Learning using AFS
Figure 13(b) shows the average NSE penalties when operating based on an NSE model learned using different querying approaches. Clusters for critical state selection were generated using KMeans clustering algorithm with K=5K\!=\!5 in the Atari Freeway domain.
Method Avg. Cost
Oracle $3759.8\pm 0.00$
Naive $61661.0\pm 0.00$
RI $71716.6\pm 0.00$
AFS (Ours) $1726.5\pm 0.00$
TABLE II: Average cost at task completion.
Table II shows the average cost for task completion. While the Naive Agent has a lower cost for task completion, it incurs the highest NSE penalty as it has no knowledge of $R_{N}$. RI causes more NSEs, as its reward function does not fully model the penalties for mild and severe NSEs. Overall, the results show that AFS consistently mitigates NSEs, without affecting the task performance substantially.
Appendix B Human Subjects Pilot Study in Simulation
Figure 14: Results from the user study on a simulated domain.
We conducted a within-subjects pilot study on a $5\!\times\!5$ Vase domain in simulation as shown in Fig. 14(a), with $12$ human participants who had completed at least one course in Reinforcement Learning. The objective of this study is to evaluate whether: (1) AFS outperforms the baselines when a feedback preference model is learned from user interactions; (2) the selected feedback formats and critical states enhance agent’s learning, and align with user preferences. The study was conducted with approval from Oregon State University IRB, and the participants were compensated with a $\mathdollar 10$ Amazon gift card for their time.
B-A Study Design
After introducing the domain and the agent’s objective, users completed a tutorial where they interacted with the system by providing feedback in each of the six formats. The study interface included feedback buttons that varied based on the format. This was followed by a calibration phase, during which the users’ preference model was learned. Each user was prompted five times per format to provide feedback, with the option to respond or ignore, allowing them to express their interaction preferences. The probability of receiving feedback in a given format was determined by the fraction of prompts the user responded to, while the cost was based on their self-reported effort.
The study comprised three phases, each evaluating a different baseline approaches to select feedback queries: (1) RI, (2) AFS with Random $\Omega$, and (3) AFS with our proposed method for critical state selection. To prevent bias, users were unaware of the approach used in each phase. After completing a phase, they were shown a trajectory of the agent’s learned policy and asked to evaluate the approach used in that phase.
B-B User Interface for Feedback Collection
Figure 15 illustrates the interface used in the simulation-based human subjects study. Participants interacted with the simulated robot through a GUI consisting of feedback buttons whose labels and available options varied depending on the feedback format. For each query, the interface would display the action the agent intends to take in the gridworld and provide a corresponding set of input buttons to record user feedback.
Figure 15: Interface for the human subjects study in simulation. Participants provide feedback via button clicks, with available options varying by format, as shown.
Approach Intelligent Feedback Critical Points (%) Improved Performance (%)
Yes No Overlap Yes No
RI $3.33\pm 1.23$ $83.30\pm 0.37$ $16.70\pm 0.37$ $73.47\pm 5.49$ $91.70\pm 0.28$ $8.30\pm 0.28$
AFS w/ Random $\Omega$ $2.82\pm 0.94$ $66.70\pm 0.47$ $33.30\pm 0.94$ $75.51\pm 5.27$ $41.70\pm 0.49$ $58.30\pm 0.49$
AFS (Ours) $3.25\pm 0.83$ $100.00\pm 0.00$ $0.00\pm 0.00$ $81.63\pm 4.94$ $100.00\pm 0.00$ $0.00\pm 0.00$
TABLE III: Participants’ qualitative assessment from the pilot study on a simulated domain.
This interface design allowed participants to provide both categorical and comparative feedback efficiently. After an initial training phase to practice providing feedback in different formats, participants self-reported the probability, $\psi(f)$, of providing feedback in a given format $f$, and effort ratings, $C(f)$.
B-C Results and Takeaways
Fig. 14(b) shows that our approach tends to result in fewer NSEs, compared to the baselines. Since the NSE penalty is an aggregate measure that obscures severity distribution, we report exact NSE encounters by category for this study. Table III reports average over responses to our questions: “On a scale of 1 to 5, how intelligent do you think the agent’s choice of feedback formats are, given your preferences?”, “Were the states in which the agent requested for feedback critical to its learning?”, and “Did the agent’s performance improve at the end of the learning phase?”. In addition, we also report the overlap between user-identified important query points and query points chosen by each approach.
Overall, the results of this pilot study indicate that (1) AFS tends to effectively select query points and lead to improved learning outcomes, when operating under a learned feedback model; and (2) AFS’s performance in this pilot study where users interact with a simulated agent is comparable to that of our results in simulation. Building on these results and the insights gained from this pilot study, we next conduct a user study where human participants interact with a physical robot. Such a setting will enable us to evaluate how well the observed trends extend to human-robot physical interactions, and how that affects the usability, trust and the users’ perceived workload when interacting with a system that learns using AFS.
Appendix C In-Person User Study with Kinova Arm
C-A User Interface and Feedback Modalities
Figure 16 illustrates the interface and feedback mechanisms used during the in-person user study with the Kinova Gen3 7-DoF robotic arm. Participants interacted with the robot through both a GUI and direct physical manipulation of the arm, depending on the feedback format being queried.
Interface Layout The GUI displayed a simulation of the robot’s motion on the tabletop, together with clearly labeled directional arrows (Up, Down, Left, Right) corresponding to the robot’s possible actions. The lower portion of the interface presented a dialog box, that indicates the current condition, the robot’s selected action, and the question or instruction corresponding to the current format. Each query consisted of a brief robot motion in simulation, followed by the corresponding GUI prompt. Participants could replay the motion before submitting their feedback.
Figure 16: Interface and prompts used in the user study with the Kinova arm. The interface displayed a short clip of the robot’s action (indicated by the white dotted arrow) along with a dialog box prompting user feedback. Top row: Approval and Correction; Bottom row: Ranking and DAM formats.